# The role of equation of state on the spin-up of millisecond pulsars

### v1

@@ Abstract
Millisecond pulsars (MSPs) are recycled pulsars which have been spun-up due to mass accretion during the phase of mass exchange in binaries. Although the interactions with companion stars play important roles on the spin-up process, the global properties of pulsars determined by the equation of state (EoS), such as mass, radius and the moment of inertia, should also play a role. We investigate the spin-up of MSPs in neutron star (NS) and strangeon star (SS) models, both of which have passed the tests by the existence of high-mass pulsars and the tidal deformability of GW 170817. Combining the spin-up condition and the transferred angular momentum, we can constrain the accreted mass and the magnetic field strength. The results indicate that it is easier for an SS to form a fully recycled low-mass MPS, whose spin-period is below 10 ms and mass is below about $1.5M_{\odot}$, than that for an NS. Finding more MSPs with low mass and and short spin-period could help to put more strict constraints on the EoS of pulsars.
dense matter — stars: neutron — pulsars: genera
@@ 1 Introduction
Considering that the EoSs should satisfy the constraints by both the existence of higher than two-solar-mass pulsars and the tidal deformability of GW 170817, we use the LX3630 EoS (Lai & Xu, 2009; Gao et al., 2022) for SSs and the AP4 EoS (Akmal & Pandharipande, 1997) for NSs. Combining the spin-up condition derived in Sec.2 and the minimal spin-period of a recycled pulsar estimated in Sec.3, we can constrain the upper limit for the magnetic field and the lower limit for the accreted mass. The results shown in Sec.4 indicate that it could be easier for an SS to form a low-mass ($M<1.5M_{\odot}$) and fully recycled ($P<10$ ms) MSP than that for an NS. Conclusions and discussions are made in Sec.5.
@@ 2 The spin-up line
The strength of the magnetic dipole field $B$ is usually estimated in terms of the spin period $P$ and its time derivative $\dot{P}$, by applying the vacuum magnetic dipole model and assuming that the rate of rotational energy loss equals the energy-loss rate caused by dipole radiation due to the inclination angle $\alpha$ (between the axis of the magnetic dipole field and the rotation axis of the pulsar). However, the real spin-down torque should be more complicated than the magnetic dipole braking in vacuum. For example, the spin-down torque could be exerted by particle wind (Li et al., 2014; Tong, 2016) or by the plasma current in the magnetosphere (Spitkovsky, 2006). To give a more general and less model-dependent investigation, in the following we keep $B$ as a parameter instead of replacing it by combination of other quantities. Then assuming the mass accretion rate $\dot{M}$ to be the Eddington accretion limit $\dot{M}_{\rm Edd}\simeq 3\times 10^{-8}M_{\odot}\rm yr^{-1}$, and using equation (1), we can plot the spin-up line in $MP$ diagram, which would give the maximum allowed magnetic field $B_{\rm max}$ for the further spin-up process.
The spin-up condition written in equation (1) is an equation connecting mass and radius, so it is EoS-dependent. Different EoS models will give different spin-up lines in $MP$ diagram. and Fig.1 shows the results for NS and SS models. Fig.1 shows the spin-up lines for both SS (red lines) and NS (blue lines) models, and the solid and dashed lines correspond to $B=5\times 10^{8}$ G and $B=2\times 10^{8}$, respectively. We can see that the spin-up lines of NSs are decreasing with the mass, and that of SSs are nearly independent of the mass. The spin-up line gives the lower limit for spin-period $P$ of recycled pulsars, so Fig.1 indicates that an SS could have smaller values of $P$ than an NS in the recycling process.
Figure 1: The spin-up lines for both SS (red lines) and NS models (blue lines). The solid and dashed lines correspond to $B=5\times 10^{8}$ G and $B=2\times 10^{8}$ G, respectively.
@@ 3 The minimal spin-period by accretion
Besides the dependence of $M$, $R$ and $I$, the minimal spin-period determined by equation (2) also depends on $r_{\rm A}$ which introduces $B$ and $\dot{M}$ as variables. To estimate the differences in the minimal spin-period of a recycled pulsar between different EoS models, we make the same assumptions as in Tauris et al. (2012) that, the magnetic field decays rapidly in the early phases of the accretion process and accreting pulsars accumulate the majority of mass while spinning near equilibrium. In addition, we take $\dot{M}=\dot{M}_{\rm Edd}$ as in Sec.2, since the dependence on $\dot{M}$ is relatively weak. Consequently, the variables in the right-hand side of equation (2) are $M_{1}$, $M_{2}$, $M$, $R$ and $I$. Given an EoS, the minimal spin-period of a recycled pulsar with the amount of accreted mass $\Delta M=M_{2}-M_{1}$ could be derived.
@@
Figure 2: The minimal spin-period by accreting the amount of mass $\Delta M$ for both SS (red lines) and NS models (blue lines), when the strength of magnetic field $B=5\times 10^{8}$ G. The solid and dashed lines correspond to $\Delta M=0.05M_{\odot}$ and $\Delta M=0.1M_{\odot}$, respectively. The horizontal ordinate denotes the final mass after accretion.
@@ 4 Comparison with observations
The recycling of pulsars is usual studied by comparison with observations in $P\dot{P}$ diagram. As we mentioned above, we keep $B$ as a parameter instead of replacing it by combination of $P$ and $\dot{P}$. Therefore, as shown in equations (1) and (2), in this paper the quantities to be compared with observations are $M$ and $P$. Because equation (2) gives the minimal value of the final spin-period $P_{\rm min}$ for a given value of $\Delta M$, for a pulsar with a known value of $P$ it gives the minimum accreted mass $\Delta M_{\rm min}$. In the $MP$ diagram, where $M$ denotes the final mass after accretion, we can plot the contours of $B_{\rm max}$ for a given EoS by combining equations (1) and (2).
In the $MP$ diagram, a higher contour of $B_{\rm max}$ corresponds to a higher value of $B_{\rm max}$, and a lower curve of $\Delta M_{\rm min}$ corresponds to a higher value of $\Delta M_{\rm min}$. Therefore, if we could determine the lower limit for $B$ and upper limit for $\Delta M$, i.e. $B\geq B_{0}$ and $\Delta M\leq\Delta M_{0}$, then the allowed region in $MP$ diagram is above both the curves of $B_{\rm max}=B_{0}$ and $\Delta M_{\rm min}=\Delta M_{0}$. Fig.3 shows the allowed regions of SSs (left) and NSs (right), where we adopt $\Delta M_{\rm min}=0.05M_{\odot}$ for solid curves and $B_{\rm max}=5\times 10^{8}$ G for dashed curves. From this example we can see that, the allowed region of SSs covers the lower left region in $MP$ diagram where NSs cannot reach, indicating that it could be easier for an SS to be fully recycled than that for an NS with the same $B$ when the mass is below about $1.5M_{\odot}$.
@@
Figure 3: The allowed regions of SSs (left) and NSs (right), where solid curves denote $\Delta M_{\rm min}=0.05M_{\odot}$ and dashed curves denote $B_{\rm max}=5\times 10^{8}$ G. The horizontal ordinate denotes the final mass after accretion. If we could determine the lower limit for the magnetic field strength $B_{0}$ and the upper limit for the amount of accreted mass $\Delta M_{0}$, then the allowed region in $MP$ diagram is above both the curves of $B_{\rm max}=B_{0}$ and $\Delta M_{\rm min}=\Delta M_{0}$.
@@
As mentioned above, to set the allowed region in $MP$ diagram we should know the the lower limit for the magnetic field strength $B_{0}$ and the upper limit for the amount of accreted mass $\Delta M_{0}$, although there are a lot of uncertainties to constrain $B$ and $\Delta M$. We could estimate $B_{0}$ by considering that, the vacuum magnetic dipole model with inclination angle $\alpha$ gives $B_{\rm dipole}\simeq 1.6\times 10^{9}\ {\rm G}\ \sqrt{P\dot{P}}\ \sin\alpha^{-1}$, so by setting $\sin\alpha=1$ we can get $B_{0}\simeq 1.6\times 10^{9}\ {\rm G}\ \sqrt{P\dot{P}}$. It is interesting to noted that, for each pulsar whose data of $M$ and $P$ is plotted in Fig.4 (see below), the value of $1.6\times 10^{9}\ {\rm G}\ \sqrt{P\dot{P}}$ calculated using the data of $P$ and $\dot{P}$ from the ATNF Pulsar Catalogue  [^1] https://www.atnf.csiro.au/research/pulsar/psrcat/  (Manchester et al., 2005) is larger than the value of $B$ of the nearest dashed curve below it in $MP$ diagram, under both SS and NS models. So setting $B_{0}$ as the lower limit of $B$ seems to be a reasonable way. It is also possible to estimate $\Delta M_{0}$ from the evolution process of the binary system, such as the accretion rate, the age of the MSP and the interaction between the MSP and the donor star.
@@
To compare with observations, we extend the allowed regions shown in Fig.3 to different values of $B_{\rm max}$ and $\Delta M_{\rm min}$, and we show in Fig.4 the results for SSs (left) and NSs (right). Solid curves correspond to $\Delta M_{\rm min}/M_{\odot}=0.02,0.04,0.06,0.1,0.15,0.3$, increasing downwards, and dashed curves correspond to contours for $B_{\rm max}/10^{8}{\rm G}=2,4,8,12$, increasing upwards. To ensure that the data are of fully recycled pulsars, we choose the masses of pulsars whose $P<10$ ms from Özel & Freire (2016) and plot their $M$ (with 1-$\sigma$ uncertainties) and $P$ in this $MP$ diagram.
@@
Figure 4: Contours of $\Delta M_{\rm min}$ and $B_{\rm max}$ for SSs (left) and NSs (right). The horizontal ordinate denotes the final mass after accretion. Solid curves correspond to $\Delta M_{\rm min}/M_{\odot}=0.02,0.04,0.06,0.1,0.15,0.3$, increasing downwards, and dashed curves correspond to contours for $B_{\rm max}/10^{8}{\rm G}=2,4,8,12$, increasing upwards. Data points with 1-$\sigma$ uncertainties are from Özel & Freire (2016).
The different effects on parameters of recycled MSPs between SS and NS models can be seen from Fig.4. Although the curves of $\Delta M_{\rm min}$ for SSs are not significantly different from that for NSs, the differences between the contours of $B_{\rm max}$ for SSs and that for NSs are visible. The dashed curves in Fig. 4 show that the magnetic field strength will play an important role in impeding the spin-up of NSs, especially for the ones with low masses, but the impeding effect would not be important to SSs. The constraints on EoS models are more significant for masses below about $1.6M_{\odot}$ than that for higher masses. For example, if the magnetic field is higher than $4\times 10^{8}$ G ($2\times 10^{8}$ G), then an NS with $M<1.6M_{\odot}$ ($M<1.2M_{\odot}$) could not be spun up to $P<3$ ms ($P<2$ ms), but an SS could. And on the other hand, if the magnetic field is higher than $8\times 10^{8}$ G, then an SS with mass approaching $2.2M_{\odot}$ could not be spun up to $P<4$ ms, but an NS could.
As more and more masses of MSPs with increasing accuracy have been being measured, we can study the recycling of pulsars by comparison with observations in $MP$ diagram. Although from the data plotted in Fig. 4 we could not yet disfavor neither SS nor NS models, finding more MSPs with $P<10$ ms with neither low mass or high mass in the future could help to put more strict constraints on the EoS of pulsars.
@@ 5 Conclusions and discussions
The global properties of pulsars determined by the EoS, such as mass, radius and the moment of inertia, could play a role in the spin-up process of accreting pulsars in binaries. We investigate the spin-up of MSPs in AP4 EoS of NS model and LX3630 EoS of the SS model. Combining the spin-up condition and the transferred angular momentum, we can constrain the accreted mass $\Delta M$ and the magnetic field strength $B$ of recycled MSPs. The allowed regions in $MP$ diagram derived by the lower limit for $B$ and the upper limit for $\Delta M$ indicate that, it could be easier for an SS to form a low-mass and fully recycled MSP than that for an NS. Although the we could not yet rule out neither SS nor NS models by the present data, finding more samples of fully recycled MSPs with neither low mass (below about $1.5M_{\odot}$) or very high mass (above about $2.2M_{\odot}$) in the future could be helpful.
@@
Although there are a lot of work about the recycling of MSPs which mainly focus on the effects of the interactions between pulsars and donor stars during binary evolution, the mass and radius are usually fixed to be typical values, such as $1.4M_{\odot}$ and 10 km. We show in this paper, however, that different global structures could bring observational differences, implying that the effects brought by EoS models should be taken into account in studying the recycling of MSPs. It is worth adding the role of EoS to other effects, such as the nature of the donor star relating to the amount of accreted mass and evolution of the magnetic field due to accretion, to study the recycling of MSPs more comprehensively.
@@
The AP4 EoS and the LX3630 EoS respectively represent EoS models of gravity-bound and self-bound, so we take them as examples to derive some universal conclusions. A gravity-bound pulsar, as shown by AP4 EoS, would have nearly unchanged radius when the mass increases from 1 to 2 $M_{\odot}$, whereas the mass of a self-bound pulsar, as shown by LX3630 EoS, is nearly proportional to the third power of radius when the mass is well below the maximum value. This difference leads to different evolutions of $r_{\rm A}$ during mass accretion, resulting in different allowed regions in $MP$ diagram of recycled pulsars. We could make a general conclusion that, it is easier for a self-bound star to form a low-mass and fully recycled MSP than that for a gravity bound star. Finding more MSPs with spin periods less than 10 milliseconds and with accurate mass-measurement in the future, e.g. by SKA and FAST, could help to put more strict constraints on the nature of pulsars.

### v2

@@ Abstract
Millisecond pulsars (MSPs) are recycled pulsars which have been spun-up due to mass accretion during the phase of mass exchange in binaries. Although the interactions with companion stars play important roles on the spin-up process, the global properties of pulsars determined by the equation of state (EoS), such as mass, radius and the moment of inertia, should also play a role. We investigate the spin-up of MSPs in neutron star (NS) and strangeon star (SS) models, both of which have passed the tests by the existence of high-mass pulsars and the tidal deformability of GW 170817. Combining the spin-up condition and the transferred angular momentum, we can constrain the accreted mass and the magnetic field strength. The magnetic field strength will be important in impeding the spin-up of NSs, especially to the ones with low masses, but the impeding effect would not be important to SSs. The results indicate that it is easier for an SS to form a fully recycled low-mass MPS, whose spin-period is below 10 ms and mass is below about $1.5M_{\odot}$, than that for an NS. Finding more low-mass and fully recycled MSPs, with accurate mass-measurement and better constraints on magnetic field strength, could help to put more strict constraints on the EoS of pulsars.
dense matter — stars: neutron — pulsars: general
@@ 1 Introduction
Considering that the EoSs should satisfy the constraints by both the existence of higher than two-solar-mass pulsars and the tidal deformability of GW 170817, we use the LX3630 EoS (Lai & Xu, 2009; Gao et al., 2022) for SSs and the AP4 EoS (Akmal & Pandharipande, 1997) for NSs. From the spin-up condition we derive the spin-up lines in $MP$ (mass vs spin-period) diagram in Sec.2, and find that under the same magnetic field strength an SS could spin faster than an NS in the recycling process. To take into account the effect of accretion process, in Sec.3 we derive the minimal spin-period by assuming that all the angular momentum of the accreted material is added the the pulsar. Although the results of Sec.3 show that NSs would spin faster than SSs with the same amount of accreted mass and the same magnetic field, the minimal-spin contours derived in Sec.4 show that to get the minimal spin the amount of accreted mass needed by an SS and an NS would be almost the same. The results shown in Sec.4 indicate that the constraints on EoS models are significant for masses below about $1.5M_{\odot}$, especially when we get more accurate estimation of the magnetic field strength. Conclusions and discussions are made in Sec.5.
@@ 2 The spin-up line
The strength of the magnetic dipole field $B$ is usually estimated in terms of the spin period $P$ and its time derivative $\dot{P}$, by applying the vacuum magnetic dipole model and assuming that the rate of rotational energy loss equals the energy-loss rate caused by dipole radiation due to the inclination angle $\alpha$ (between the axis of the magnetic dipole field and the rotation axis of the pulsar). However, the real spin-down torque should be more complicated than the magnetic dipole braking in vacuum. For example, the spin-down torque could be exerted by particle wind (Li et al., 2014; Tong, 2016) or by the plasma current in the magnetosphere (Spitkovsky, 2006). To give a more general and less model-dependent investigation, in the following we keep $B$ as a parameter instead of replacing it by combination of other quantities. Then assuming the mass accretion rate $\dot{M}$ to be the Eddington accretion limit $\dot{M}_{\rm Edd}\simeq 3\times 10^{-8}M_{\odot}\rm yr^{-1}$, and using equation (1), we can plot the spin-up line in $MP$ diagram.
The spin-up condition written in equation (1) is an equation connecting mass and radius, so it is EoS-dependent. Different EoS models will give different spin-up lines in $MP$ diagram, and Fig.1 shows the results for NS and SS models. Fig.1 shows the spin-up lines for both SS (red lines) and NS (blue lines) models, and the solid and dashed lines correspond to $B=5\times 10^{8}$ G and $B=2\times 10^{8}$, respectively. We can see that the spin-up lines of NSs are decreasing with the mass, and that of SSs are nearly independent of the mass. The spin-up line gives the lower limit for spin-period $P$ of recycled pulsars, so Fig.1 indicates that an SS could have smaller values of $P$ than an NS in the recycling process.
Figure 1: Spin-up lines for both SS (red lines) and NS models (blue lines). The solid and dashed lines correspond to $B=5\times 10^{8}$ G and $B=2\times 10^{8}$ G, respectively.
@@ 3 The minimal spin-period by accretion
Besides the dependence of $M$, $R$ and $I$, the minimal spin-period determined by equation (2) also depends on $r_{\rm A}$ which introduces $B$ and $\dot{M}$ as variables. To estimate the differences in the minimal spin-period of a recycled pulsar between different EoS models, we make the same assumption as in Tauris et al. (2012) that, the magnetic field decays rapidly in the early phases of the accretion process and accreting pulsars accumulate the majority of mass while spinning near equilibrium (but this assumption can be relaxed in deriving the minimal-spin contour which contains different values of $B$, see Sec.4). In addition, we take $\dot{M}=\dot{M}_{\rm Edd}$ as in Sec.2, since the dependence on $\dot{M}$ is relatively weak. Consequently, the variables in the right-hand side of equation (2) are $M_{1}$, $M_{2}$, $M$, $R$ and $I$. Given an EoS, the minimal spin-period of a recycled pulsar with the amount of accreted mass $\Delta M=M_{2}-M_{1}$ could be derived.
@@
Figure 2: The minimal spin-period by accreting the amount of mass $\Delta M$ for both SS (red lines) and NS models (blue lines), when the strength of magnetic field $B=5\times 10^{8}$ G. The solid and dashed lines correspond to $\Delta M=0.05M_{\odot}$ and $\Delta M=0.1M_{\odot}$, respectively. In $MP$ diagram, the horizontal ordinate denotes the final mass after accretion.
@@ 4 Comparison with observations
The recycling of pulsars is usual studied by comparison with observations in $P\dot{P}$ diagram. As we mentioned above, we keep $B$ as a parameter instead of replacing it by combination of $P$ and $\dot{P}$. Therefore, as shown in equations (1) and (2), in this paper the quantities to be compared with observations are $M$ and $P$.
If the values of both $B$ and $\Delta M$ of a recycled pulsar is known, we can plot the allowed region in $MP$ diagram which satisfies both equations (1) and (2), i.e. by putting Fig.1 and Fig.2 together. We show the allowed regions of SSs (left) and NSs (right) for $\Delta M=0.05M_{\odot}$ and $B=5\times 10^{8}$ G in Fig.3, where dashed line denotes the general spin-up line (cf. Fig.1) and solid line shows the minimal spin-period by accretion (cf. Fig.2). From this example we can see that, the allowed region of SSs covers the lower left region in $MP$ diagram where NSs cannot reach, indicating that it could be easier for an SS to be fully recycled than that for an NS with the same $B$ when the mass is below about $1.5M_{\odot}$.
@@
Figure 3: The allowed regions of SSs (left) and NSs (right) in $MP$ diagram, where solid curves denote minimal-spin lines with $\Delta M=0.05M_{\odot}$ and $B=5\times 10^{8}$ G, and dashed curves denote spin-up lines with $B=5\times 10^{8}$ G.
@@
However, in most cases we are not certain about both the value of $B$ and $\Delta M$ for recycled pulsars. If we estimate $B$ by considering that, the vacuum magnetic dipole model with inclination angle $\alpha$ gives $B_{\rm dipole}\simeq 1.6\times 10^{9}\ {\rm G}\ \sqrt{P\dot{P}}\ \sin\alpha^{-1}$, then by setting $\sin\alpha=1$ we can get $B>1.6\times 10^{9}\ {\rm G}\ \sqrt{P\dot{P}}$. It is interesting to noted that, each pulsar whose data of $M$ and $P$ is plotted in Fig.5 (see below) is above the spin-up line of $B=1.6\times 10^{9}\ {\rm G}\ \sqrt{P\dot{P}}$ calculated using the data of $P$ and $\dot{P}$ from the ATNF Pulsar Catalogue (https://www.atnf.csiro.au/research/pulsar/psrcat/(Manchester et al., 2005) under both SS and NS models. As mentioned before, in this paper we keep $B$ as a parameter, instead of replacing in some model-dependent ways. Then how to show the difference between SSs and NSs models by comparing with observations?
Combining equations (1) and (2) we can get the minimal-spin contour by accretion of $\Delta M$, which is defined as the contour of $\Delta M$ in $MP$ diagram, representing the minimal spin-period by accretion of $\Delta M$ for different values of $B$. For a data point in $MP$ diagram, there exists both a minimal-spin contour and a spin-up line intersecting it. The value of $\Delta M$ of the minimal-spin contour is the minimal accreted mass of the pulsar, while the value of $B$ of the spin-up line is the upper limit on its magnetic field strength. If we can use some other methods to constrain $\Delta M$ (e.g. from the binary evolution) and $B$ (e.g. theoretically from the magnetic field decay in accretion, or observationally from the cyclotron absorption lines), it would be hopeful to see that whether this pulsar is more likely to be an NS or an SS.
From Fig.3 we can infer that low mass and fast spinning pulsars would present significant different parameters between SS and NS models, so we use a hypothetical data point in $MP$ diagram as an example in Fig.4. If this pulsar is an SS, a minimal-spin contour of $\Delta M=0.059M_{\odot}$ (red line) and a spin-up line of $B=6.6\times 10^{8}$ G (red dashed line) intersect it in $MP$ diagram, indicating that the accreted mass $\Delta M>0.059M_{\odot}$ and its magnetic field strength $B<6.6\times 10^{8}$ G. One the other hand, if this pulsar is an NS, a minimal-spin contour of $\Delta M=0.060M_{\odot}$ (blue line) and a spin-up line of $B=4.1\times 10^{8}$ G (blue dashed line) intersect it in $MP$ diagram, indicating that the accreted mass $\Delta M>0.060M_{\odot}$ and its magnetic field strength $B<4.1\times 10^{8}$ G. This example shows that the difference in the the upper limit on $B$ between SS and NS models is significant for low mass and fast spinning pulsars.
Figure 4: We use a hypothetical data point in $MP$ diagram as an example. If this pulsar is an SS, a minimal-spin contour of $\Delta M=0.059M_{\odot}$ (red line) and a spin-up line of $B=6.6\times 10^{8}$ G (red dashed line) intersect it, indicating that the accreted mass $\Delta M>0.059M_{\odot}$ and its magnetic field strength $B<6.6\times 10^{8}$ G. If this pulsar is an NS, a minimal-spin contour of $\Delta M=0.060M_{\odot}$ (blue line) and a spin-up line of $B=4.1\times 10^{8}$ G (blue dashed line) intersect it, indicating that the accreted mass $\Delta M>0.060M_{\odot}$ and its magnetic field strength $B<4.1\times 10^{8}$ G.
@@
To compare with actual data, we extend the curves in Fig.4 to different values of $B$ and $\Delta M$, and we show in Fig.5 the results for SSs (left) and NSs (right). Solid curves correspond to minimal-spin contours with $\Delta M/M_{\odot}=0.02,0.04,0.06,0.1,0.15,0.3$, increasing downwards, and dashed curves correspond to spin-up lines with $B_{\rm max}/10^{8}{\rm G}=2,4,8,12$, increasing upwards. To ensure that the data are of fully recycled pulsars, we choose the masses of pulsars whose $P<10$ ms from Özel & Freire (2016) and plot their $M$ (with 1-$\sigma$ uncertainties) and $P$ in this $MP$ diagram.
@@
Figure 5: Minimal-spin contours of $\Delta M$ and spin-up lines of $B$ for SSs (left) and NSs (right). The horizontal ordinate denotes the final mass after accretion. Solid curves correspond to $\Delta M/M_{\odot}=0.02,0.04,0.06,0.1,0.15,0.3$, increasing downwards, and dashed curves correspond to contours for $B/10^{8}{\rm G}=2,4,8,12$, increasing upwards. Data points with 1-$\sigma$ uncertainties are from Özel & Freire (2016).
The different effects on parameters of recycled MSPs between SS and NS models can be seen from Fig.5. Although the minimal-spin contours of $\Delta M$ for SSs are not significantly different from that for NSs, the differences between the contours of $B$, i.e. the spin-up lines, for SSs and that for NSs are visible. The dashed curves in Fig. 5 show that the magnetic field strength will play an important role in impeding the spin-up of NSs, especially for the ones with low masses, but the impeding effect would not be important to SSs. The constraints on EoS models are more significant for masses below about $1.5M_{\odot}$ than that for higher masses.
If we can find some fast spinning pulsars with low masses, and if we can additionally know the lower limit of $B$, i.e. $B>B_{0}$, it would be hopeful to see that whether these pulsars is more likely to be an NS or an SS. For example, if the magnetic field is higher than $4\times 10^{8}$ G ($2\times 10^{8}$ G), then an NS with $M<1.6M_{\odot}$ ($M<1.2M_{\odot}$) could not be spun up to $P<3$ ms ($P<2$ ms), but an SS could. As we mentioned before, each pulsar plotted in Fig.5 is above its spin-up line of $B=1.6\times 10^{9}\ {\rm G}\ \sqrt{P\dot{P}}$ calculated using its data of $P$ and $\dot{P}$, so at this point if we take the lower limit magnetic field strength $B_{0}$ as $1.6\times 10^{9}\ {\rm G}\ \sqrt{P\dot{P}}$ we could not yet disfavor neither SS nor NS models. Certainly, finding more MSPs of $P<10$ ms with low mass and more accurate estimation of $B_{0}$ in the future could help to put more strict constraints on the EoS of pulsars.
@@ 5 Conclusions and discussions
The global properties of pulsars determined by the EoS, such as mass, radius and the moment of inertia, could play a role in the spin-up process of accreting pulsars in binaries. We investigate the spin-up of MSPs in AP4 EoS of NS model and LX3630 EoS of the SS model. Combining the spin-up condition and the transferred angular momentum, we can constrain the accreted mass $\Delta M$ and the magnetic field strength $B$ of recycled MSPs. Under the same magnetic field strength, an SS could spin faster than an NS in the recycling process, and NSs would spin faster than SSs with the same amount of accreted mass. The minimal-spin contours show that to get the minimal spin the amount of accreted mass needed by an SS and an NS would be almost the same, but the differences between the spin-up lines for SSs and that for NSs are obvious. The magnetic field strength will play an important role in impeding the spin-up of NSs, especially for the ones with low masses, but the impeding effect would not be important to SSs, so it could be easier for an SS to form a low-mass and fully recycled MSP than that for an NS.
The constraints on EoS models are more significant for masses below about $1.5M_{\odot}$, when the upper limit of the allowed magnetic field strength of an SS is higher than that of an NS. To give a more general and model-independent investigation, in this paper we keep the magnetic field strength $B$ as a parameter instead of replacing it in some model-dependent ways. It is worth noting that an accreting pulsar would have an lower limit of magnetic field strength (e.g. see Zhang & Kojima (2006)), so the constraints on NS model would be more strict than that on SS model. For the fast spinning pulsars with low masses, if we can additionally derive the lower limit of magnetic field strength theoretically from the magnetic field decay models, or if we can directly detect the magnetic field strength from the cyclotron absorption lines, it would be hopeful to see that whether these pulsars is more likely to be an NS or an SS.
@@
Although there are a lot of work about the recycling of MSPs which mainly focus on the effects of the interactions between pulsars and donor stars during binary evolution, the mass and radius are usually fixed to be typical values, such as $1.4M_{\odot}$ and 10 km. We show in this paper, however, that different global structures could bring observational differences, implying that the effects brought by EoS models should be taken into account in studying the recycling of MSPs. To focus on the EoS-dependent results, it is inevitable to ignore the dependence on other factors in this paper, such as the pattern of disk-fed accretion flow and the magnetic field configuration outside the accreting star, which lead some simplifications in equations (1) and (2). In addition we do not consider the nature of the donor star relating to the amount of accreted mass and evolution of the magnetic field due to accretion. Still, our results indicate that it is worth adding the role of EoS to other effects to study the recycling of MSPs more comprehensively.
@@
The AP4 EoS and the LX3630 EoS respectively represent EoS models of gravity-bound and self-bound, so we take them as examples to derive some universal conclusions. A gravity-bound pulsar, as shown by AP4 EoS, would have nearly unchanged radius when the mass increases from 1 to 2 $M_{\odot}$, whereas the mass of a self-bound pulsar, as shown by LX3630 EoS, is nearly proportional to the third power of radius when the mass is well below the maximum value. This difference leads to different evolutions of $r_{\rm A}$ during mass accretion, resulting in different allowed regions in $MP$ diagram of recycled pulsars. We could make a general conclusion that, it is easier for a self-bound star to form a low-mass and fully recycled MSP than that for a gravity bound star. Although the we could not yet rule out neither SS nor NS models by the present data, finding more samples of fully recycled low-mass MSPs with accurate mass-measurement and better constraints on magnetic field strength in the future, e.g. by SKA and FAST, could be helpful to put more strict constraints on the nature of pulsars.